OpenStudy (anonymous):

Find the value of k so that (–1, 5) lies on kx – 4y – k = 0. A. 20 B. –10 C. infinite number of solutions D. no solution

OpenStudy (anonymous):

Can anyone explain how to solve this?

ganeshie8 (ganeshie8):

plugin x = -1, and y = 5

ganeshie8 (ganeshie8):

in the given equation

ganeshie8 (ganeshie8):

kx – 4y – k = 0 ^ ^

ganeshie8 (ganeshie8):

k(-1) - 4(5) - k = 0

ganeshie8 (ganeshie8):

solve k

OpenStudy (anonymous):

Does k equal -20?

OpenStudy (anonymous):

k(-1) -20=0, k(-1)=20, k=-20

OpenStudy (anonymous):

Is this correct?

ganeshie8 (ganeshie8):

you forgot the -k in equation. do it agian :)

OpenStudy (anonymous):

k=-10.....

OpenStudy (anonymous):

-k-20-k=...-2k-20=...-2k=20=.....k=-10

OpenStudy (anonymous):

k(-1)-k = 20. How do I continue from here?

OpenStudy (anonymous):

-k-k=20........-2k=20...........k=-10

OpenStudy (anonymous):

Thank you soooo much :)

ganeshie8 (ganeshie8):

yw :)